OpenStudy (anonymous):

Factor the trinomial: x2 - 10x + 9 (x - 1)(x - 9) (x + 1)(x - 10) (x - 3)(x - 3) prime

OpenStudy (callisto):

x^2 - 10x + 9 =x^2 - 9x -x + 9 =x(x-9) -(x-9) Can you try to group the like terms and see what you'll get?

OpenStudy (anonymous):

I don't know what 'group these terms' means.

OpenStudy (callisto):

Hmmm... In x(x-9) and -(x-9) what is the common factor?

OpenStudy (anonymous):

9?

OpenStudy (callisto):

Nope Take 6 and 9 as an example 6 = 2(3), 9=3(3) The common factor is 3 and you see they are both multiples of 3 Can you try again now?

OpenStudy (anonymous):

Still doesn't make any sense.

OpenStudy (callisto):

which part does not make sense?

OpenStudy (anonymous):

all of it

OpenStudy (callisto):

how do you factorise xy +x ?

OpenStudy (anonymous):

No idea

OpenStudy (callisto):

Do you know what is common factor?

OpenStudy (anonymous):

no

OpenStudy (callisto):

Alright, then please visit this website to get the basic idea http://www.purplemath.com/modules/factnumb.htm

OpenStudy (callisto):

Tell me if you get the basic idea of it after reading the webpage. If you still don't understand, I'll try to explain it to you again.

OpenStudy (callisto):

After you've finished reading and know the basic of factorization, we can start discussing your problem

OpenStudy (callisto):

@BazBendell I don't feel safe to ask you read the webpage.... Sigh, we shall start from the basic here... First you need to know that 6 = 1x6 = 2x3, understand this?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Okay, so ,1, 2, 3, and 6 would be the factor of 6, got this part?

OpenStudy (anonymous):

yup

OpenStudy (callisto):

for 9 , it can be expressed as 1x9, and 3x3 So, what is the factors of 9?

OpenStudy (callisto):

*is -> are

OpenStudy (anonymous):

1 and 3

OpenStudy (callisto):

and also?

OpenStudy (anonymous):

9?

OpenStudy (callisto):

yes

OpenStudy (callisto):

so factors of 6 are 1 2 3 and 6 factors of 9 are 1 3 and 9 What is the common factor of 6 and 9?

OpenStudy (callisto):

@BazBendell

OpenStudy (anonymous):

3

OpenStudy (callisto):

yes, so do you understand what common factor is?

OpenStudy (anonymous):

for the most part.

OpenStudy (callisto):

So what are the common factors of 2xy and 6x^2y^2?

OpenStudy (callisto):

@BazBendell

OpenStudy (callisto):

Hint: 2xy = 2 (x) (y) 6x^2y^2 = (2)(3)(x)(x)(y)(y) What are the common factors?